# Fast Frank–Wolfe Algorithms with Adaptive Bregman Step-Size for Weakly Convex Functions

### v3

@@ 2 Preliminaries
The HEB condition shows sharpness bounds on the primal gap (see [bolte17, alex17sharp]), which has also been extensively analyzed for FW algorithms [KDP2018, KDP2018jour]. Convergence for nonconvex optimization under the quadratic growth or HEB condition was established in [Davis2024-rm, Davis2018-bb, Davis2024-og]. We employ the HEB condition for convex optimization, while we use the local quadratic growth for nonconvex optimization.
@@ 3 Proposed Bregman FW Algorithms
  1. (i)
    $\phi\in\mathcal{G}(C)$ with $\operatorname{cl}{C}=\operatorname{cl}{\operatorname{dom}\phi}$ is strictly convex on $C\equiv\operatorname{int}\operatorname{dom}\phi$;
  2. (ii)
    $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C$;
  3. (iii)
    The pair $(f,\phi)$ is $L$-smad on $P$;
  4. (iv)
    $P\subset\mathbb{R}^{n}$ is a nonempty compact convex set with $P\subset C$.
Assumption 3.1(i)-(iii) are standard for Bregman-type algorithms [Bolte2018-zt, Takahashi2024-wz]. The strict convexity of $\phi$ ensures the existence of $\nu$ by Lemma A.3 and is satisfied by many kernel generating distances (see, e.g., [Bauschke2017-hg, Bauschke1997-vk], and [Dhillon2008-zz, Table 2.1]). Assumption 3.1(iv) is standard for FW algorithms, and $P\subset C$ is natural ensuring that $D_{\phi}$ is well-defined on $P$. The original FW algorithm is given in Algorithm 1.
@@
3      $v_{t}\leftarrow\operatorname*{argmin}_{v\in P}\langle\nabla f(x_{t}),v\rangle$
4     $x_{t+1}\leftarrow(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$
@@
In what follows, let $x^{*}\in\operatorname*{argmin}_{x\in P}f(x)$ and $f^{*}=f(x^{*})$. We recall a key property for analyzing convergence rates in convex optimization and derive its counterpart in weakly convex optimization.
Lemma 3.2 (Primal gap, dual gap, and Frank–Wolfe gap for convex $f$ [Pokutta2024-qu, Lemma 4.1]).
$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle.$ (3.1)
Lemma 3.3 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex $f$).
$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (3.2)
@@ 3.1 Bregman Step-Size Strategy for $\gamma_{t}$
Assume that $(f,\phi)$ is $L$-smad. Using the optimality condition of the right-hand side of (2.2) in terms of $\gamma$, we find
@@
Theoretically, when $x,v\in P$ and $\gamma\in[0,1]$, the Frank–Wolfe step $x_{+}=(1-\gamma)x+\gamma v\in P$ is well-defined. We update $\gamma=\min\left\{\left(\frac{\langle\nabla f(x),x-v\rangle}{L(1+\nu)D_{\phi}(v,x)}\right)^{1/\nu},\gamma_{\max}\right\}$ with some $\gamma_{\max}\in\mathbb{R}$ (usually set $\gamma_{\max}=1$). This step-size strategy provides the progress lemma.
Lemma 3.4 (Progress lemma from the Bregman short step-size).
Suppose that Assumption 3.1 holds. Define $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$. Then if $x_{t+1}\in\operatorname{dom}f$, $f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ for $0\leq\gamma_{t}\leq\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu}$.
@@
2      Choose $\beta<1$, $\eta\leq 1$, and $\tau>1$
3      $M\leftarrow\eta\tilde{L}$, $\kappa\leftarrow 1$
4      while $D_{f}(x_{+},x)\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ do
@@ 6           $x_{+}\leftarrow(1-\gamma)x+\gamma v$
7           $M\leftarrow\tau M$
8           $\kappa\leftarrow\beta\kappa$
9
@@ 10           $\tilde{L}^{*}\leftarrow M$, $\tilde{\nu}^{*}\leftarrow\kappa$
The exact value or the tight upper bound of $L$ is often unknown, however, the algorithm’s performance heavily depends on it; an underestimation of $L$ might lead to non-convergence, and an overestimation of $L$ might lead to slow convergence. Moreover, a worst-case $L$ might be too conservative for regimes where the function is better behaved. Due to all these reasons, Pedregosa et al. [Pedregosa2020-ay] proposed an adaptive step-size strategy for FW algorithms in the case where $\phi=\frac{1}{2}\|\cdot\|^{2}$. We present our algorithm in Algorithm 2. This step-size strategy can be used as a drop-in replacement. For example, inserting $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},1)$ between lines 1 and 1 in Algorithm 1, we obtain that Algorithm 2 searches $L$ and $\nu$ satisfying (2.2).
@@
By the $L$-smad property of $(f,\phi)$ and Lemma 2.4, we know that (2.2) holds for all $M\geq L$ and $\nu\geq\kappa>0$. Therefore, the loop in line 2 in Algorithm 2 is well-defined and guaranteed to terminate. It suffices to update $\kappa$ in line 2 only if (2.1) does not hold. Thus, $\kappa$ cannot become too small.
@@ 1 $\mathcal{S}_{0}\leftarrow\{x_{0}\}$, $\lambda_{x_{0},0}\leftarrow 1$
4      $v_{t}^{\textup{FW}}\leftarrow\operatorname*{argmin}_{v\in P}\langle\nabla f(x_{t}),v\rangle$, $v_{t}^{\textup{A}}\leftarrow\operatorname*{argmax}_{v\in\mathcal{S}_{t}}\langle\nabla f(x_{t}),v\rangle$
@@
In the case where $P$ is a polytope, the classical FW algorithm might zigzag when approaching the optimal face and, in consequence, converge slowly. Away-step FW algorithm overcomes this drawback. Away steps allow the algorithm to move away from vertices in a convex combination of $x_{t}$, which can effectively shortcircuit the zigzagging. The convergence properties of this algorithm were unknown for a long time, and it was only quite recently that Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] established the linear convergence of the away-step FW algorithm. Inspired by [GueLat1986-bg, Lacoste-Julien2015-gb, Wolfe1970-ax], we propose a variant of the away-step FW algorithm utilizing Bregman distances as given in Algorithm 3.
For the following discussion, we introduce some notions. In the same way [Braun2022-ew], we define an active set $\mathcal{S}\subset\operatorname{Vert}P$, where $\operatorname{Vert}P$ denotes the set of vertices of $P$, and the away vertex as $v_{t}^{\textup{A}}\in\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x_{t}),v\rangle$, where $x_{t}$ is a (strict) convex combination of elements in $\mathcal{S}$, i.e., $x_{t}=\sum_{v\in\mathcal{S}}\lambda_{v}x_{t}$ with $\lambda_{v}>0$ for all $v\in\mathcal{S}$ and $\sum_{v\in\mathcal{S}}\lambda_{v}=1$. If $\nu$ is unknown, we can add line search procedures to search $\nu$ until (2.1) holds. If $\nu$ and $L$ are unknown, we can use $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$ with $\gamma_{t}\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{M(1+\kappa)D_{\phi}(v_{t},x_{t})}\right)^{1/\kappa},\gamma_{t,\max}\right\}$ instead of line 2 in Algorithm 2. When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu\equiv 1$, Algorithm 3 corresponds to the Euclidean away-step FW algorithm.
@@ 4 Linear Convergence for Convex Optimization
We will now establish faster convergence rates than $O(1/t)$ up to linear convergence depending on the choice of parameters. First, we establish the faster convergence of Algorithm 1 with the Bregman short step-size, i.e., $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$, in the case where the optimal solution lies in the relative interior. Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$.
@@
Suppose that Assumptions 3.1 and 4.1 hold. Let $\nu>0$ and let $f$ satisfy the HEB condition (2.3) with $q\geq 1+\nu$ and $\mu>0$ and $D$ be the diameter of $P$. Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$. Consider the iterates of Algorithm 1 with short step-size $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$ or the adaptive Bregman step-size strategy (Algorithm 2). Then, it holds that, for all $t\geq 1$,
@@
Next, we establish the linear convergence of Algorithm 3. Recall that the pyramidal width is the minimal $\delta>0$ satisfying Lemma B.9 (see [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3] for an in-depth discussion). Under Assumption 4.3, it holds that $\delta\leq D$, and this will be important in establishing convergence rates of Algorithm 3. We make the following assumption:
@@ 5 Local Linear Convergence for Nonconvex Optimization
In this section, we consider a nonconvex objective function. We establish global sublinear, local sublinear, and local linear convergence (see, for global and local sublinear, Appendices D.1 and D.2). We show that the FW algorithm converges to a minimizer $x^{*}$ when an initial point is close enough to $x^{*}$. We need weak convexity and local quadratic growth for local convergence.
@@
$f$ is $\rho$-weakly convex and local $\mu$-quadratic growth with $\zeta>0$.
@@
Finally, we establish the local linear convergence of Algorithm 3. We also assume $\rho\leq L$, which is not restrictive (see also the discussion in Appendix D.2 and Example D.4).
@@
Without loss of generality, we set $\sigma=1$. If $\rho\leq L$ does not hold, we can use $f(x_{1})-f^{*}\leq(\rho+L)D^{2}$, instead of $f(x_{1})-f^{*}\leq 2LD^{2}$ as the initial bound. Moreover, when $P$ is $\alpha$-strongly convex set and $\phi=\frac{1}{2}\|\cdot\|^{2}$, we also establish local linear convergence. Its proof can be found in Appendix D.5.

### v4

@@ 2 Preliminaries
For example, $f(x)=\log(1+x^{2})$ is nonconvex but satisfies local quadratic growth. This function is used for image restoration (Boţ et al., 2016; Stella et al., 2017). See, for more examples, Examples E.4 and E.5. The HEB condition shows sharpness bounds on the primal gap (see Bolte et al. (2017); Roulet and d’Aspremont (2017)), which has been extensively analyzed for FW algorithms (Kerdreux et al., 2019; 2022). Convergence of (sub)gradient algorithms for nonconvex optimization under the quadratic growth or HEB condition was established in Davis et al. (2018; 2024); Davis and Jiang (2024). We assume the HEB condition for convex optimization and the local quadratic growth for nonconvex optimization.
@@ 3 Proposed Bregman FW Algorithms
(i) $\phi\in\mathcal{G}(C)$with $\operatorname{cl}{C}=\operatorname{cl}{\operatorname{dom}\phi}$ is strictly convex on $C\equiv\operatorname{int}\operatorname{dom}\phi$ and has $\nu>0$ satisfying Equation 2.1; (ii) $f:\mathbb{R}^{n}\to(-\infty,+\infty]$is proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C$; (iii) The pair $(f,\phi)$ is $L$-smad on $P$; (iv) $P\subset\mathbb{R}^{n}$is a nonempty compact convex set with $P\subset C$.
@@
[…]
@@
Assumption 3.1(i)-(iii) are standard for Bregman-type algorithms (Bolte et al., 2018; Takahashi and Takeda, 2025). The strict convexity of $\phi$ ensures the existence of $\nu$ by Lemma B.3 and is satisfied by many kernel generating distances (see, e.g., Bauschke et al. (2017); Bauschke and Borwein (1997), and (Dhillon and Tropp, 2008, Table 2.1)). Assumption 3.1(iv) is standard for FW algorithms, and $P\subset C$ naturally ensures that $D_{\phi}$ is well-defined on $P$. The original FW algorithm is given in Algorithm 1. In what follows, let $x^{*}\in\operatorname*{arg\,min}_{x\in P}f(x)$ and $f^{*}=f(x^{*})$.
@@ 3.1 Bregman Step-Size Strategy for $\gamma_{t}$
Assume that $(f,\phi)$ is $L$-smad. Maximizing the right-hand side of Equation 2.2 in terms of $\gamma$, we find
@@
Theoretically, when $x,v\in P$ and $\gamma\in[0,1]$, the Frank–Wolfe step $x_{+}=(1-\gamma)x+\gamma v\in P$ is well-defined. We update $\gamma=\min\left\{\left(\frac{\langle\nabla f(x),x-v\rangle}{L(1+\nu)D_{\phi}(v,x)}\right)^{1/\nu},\gamma_{\max}\right\}$ with some $\gamma_{\max}\in\mathbb{R}$ (usually set $\gamma_{\max}=1$). This step-size strategy provides $f(x)-f(x_{+})\geq\frac{\nu}{1+\nu}\gamma\langle\nabla f(x),x-v\rangle$ for $0\leq\gamma\leq\left(\frac{\langle\nabla f(x),x-v\rangle}{L(1+\nu)D_{\phi}(v,x)}\right)^{1/\nu}$. Its proof can be found in Lemma B.8.
@@
1 Procedure $\textnormal{{step\_size}}(f,\phi,x,v,\tilde{L},\gamma_{\max})$
2  Choose $\beta\in(0,1)$, $\eta\in(0,1]$, and $\tau>1$
@@
4loop
@@ 6   $x_{+}\leftarrow(1-\gamma)x+\gamma v$
7   if $D_{f}(x_{+},x)\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ then
@@ 11    $M\leftarrow\tau M$, $\kappa\leftarrow\beta\kappa$
The exact value or the tight upper bound of $L$ is often unknown; however, the algorithm’s performance heavily depends on it; an underestimation of $L$ might lead to non-convergence, and an overestimation of $L$ might lead to slow convergence. Moreover, a worst-case $L$ might be too conservative for regimes where the function is better behaved. For these reasons, Pedregosa et al. (2020) proposed an adaptive step-size strategy for FW algorithms in the case where $\phi=\frac{1}{2}\|\cdot\|^{2}$. We present our algorithm in Algorithm 2. This step-size strategy can be used as a drop-in replacement. For example, inserting $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},1)$ between lines 1 and 1 in Algorithm 1, we obtain that Algorithm 2 searches $L$ and $\nu$ satisfying Equation 2.2. For $\phi=\frac{1}{2}\|\cdot\|^{2}$, Algorithm 2 corresponds to a (Euclidean) adaptive step-size strategy (Pedregosa et al., 2020).
@@
By the $L$-smad property of $(f,\phi)$ and Lemma B.7, we know that Equation 2.2 holds for all $M\geq L$ and $\nu\geq\kappa>0$. Therefore, the condition in line 2 in Algorithm 2 is well-defined and guaranteed to terminate. It suffices to update $\kappa$ in line 2 only if Equation 2.1 does not hold. Thus, $\kappa$ cannot become too small.
We have a bound on the total number of evaluations of Algorithm 2. Its proof can be found in Appendix B.3.
Theorem 3.3.
Let $L_{-1}$ be the initial $L$-smad estimate and $n_{t}$ be the total number of evaluations of Equation 2.2 up to iteration $t$. Then we have $n_{t}\leq\max\{(1-\log\eta/\log\tau)(t+1)+\max\{\log(\tau L/L_{-1}),0\}/\log\tau,(1+\log\nu/\log\beta)(t+1)\}$.
Pedregosa et al. (2020) showed that, asymptotically, no more than 16% of the iterations require more than one evaluation of the line search procedures when $\eta=0.9$ and $\tau=2$. This follows from the bound $(1-\log\eta/\log\tau)\leq 1.16$.
@@ 3.2 Bregman Away-Step Frank–Wolfe Algorithm
In the case where $P$ is a polytope, the classical FW algorithm might zigzag when approaching the optimal face and, in consequence, converge slowly. The away-step FW algorithm overcomes this drawback. Away steps allow the algorithm to move away from vertices in a convex combination of $x_{t}$, which can effectively short-circuit the zigzagging. The convergence properties of this algorithm were unknown for a long time, and it was only quite recently that Lacoste-Julien and Jaggi (2015) established the linear convergence of the away-step FW algorithm. Inspired by GuéLat and Marcotte (1986); Lacoste-Julien and Jaggi (2015); Wolfe (1970), we propose a variant of the away-step FW algorithm utilizing Bregman distances as given in Algorithm 3. The main difference between Algorithm 3 and the existing away-step FW algorithm lies in the update in line 3.
For the following discussion, we introduce some notions. In the same way as Braun et al. (2025), we define an active set $\mathcal{S}\subset\operatorname{Vert}P$, where $\operatorname{Vert}P$ denotes the set of vertices of $P$, and the away vertex as $v_{t}^{\textup{A}}\in\operatorname*{arg\,max}_{v\in\mathcal{S}}\langle\nabla f(x_{t}),v\rangle$, where $x_{t}$ is a (strict) convex combination of elements in $\mathcal{S}$, i.e., $x_{t}=\sum_{v\in\mathcal{S}}\lambda_{v}v$ with $\lambda_{v}>0$ for all $v\in\mathcal{S}$ and $\sum_{v\in\mathcal{S}}\lambda_{v}=1$. If $\nu$ is unknown, we can add line search procedures to search $\nu$ until Equation 2.1 holds. If $\nu$ and $L$ are unknown, we can use $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$ with $\gamma_{t}\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{M(1+\kappa)D_{\phi}(v_{t},x_{t})}\right)^{1/\kappa},\gamma_{t,\max}\right\}$ instead of line 2 in Algorithm 2. When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu\equiv 1$, Algorithm 3 corresponds to the Euclidean away-step FW algorithm.
@@ 1 $\mathcal{S}_{0}\leftarrow\{x_{0}\}$, $\lambda_{x_{0},0}\leftarrow 1$
[…]
@@
6if $\langle\nabla f(x_{t}),x_{t}-v_{t}^{\textup{FW}}\rangle\geq\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-x_{t}\rangle$ then
@@ 4 Linear Convergence for Convex Optimization
Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$. We will now establish faster convergence rates than $O(1/t)$ up to linear convergence depending on $\nu$ in Equation B.2 and $q$ in Definition 2.4. First, we establish the faster convergence of Algorithm 1 with the Bregman short step-size, i.e., $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$, or Algorithm 2 in the case where the optimal solution lies in the relative interior. Its proof can be found in Appendix D.2.
@@
Suppose that Assumptions 3.1 and 4.1 hold. Let $\nu>0$ and let $f$ satisfy the HEB condition with $q\geq 1+\nu$ and $\mu>0$ and $D$ be the diameter of $P$. Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$. Consider the iterates of Algorithm 1 with short step-size $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$ or the adaptive Bregman step-size strategy (Algorithm 2). Then, it holds that, for all $t\geq 1$,
@@
Next, we establish the linear convergence of Algorithm 3. Recall that the pyramidal width is the minimal $\delta>0$ satisfying Lemma C.9 (see (Braun et al., 2025, Lemma 2.26) or (Lacoste-Julien and Jaggi, 2015, Theorem 3) for an in-depth discussion). Under Assumption 4.3, it holds that $\delta\leq D$, and this will be important in establishing convergence rates of Algorithm 3. We make the following assumption:
@@
We have the linear convergence of Algorithm 3. Its proof can be found in Appendix D.3.
@@
When $\phi=\frac{1}{2}\|\cdot\|^{2}$, i.e., $\nu=1$, Theorems 4.2 and 4.4 correspond to existing results (Braun et al., 2025; Kerdreux et al., 2022).
@@ 5 Local Linear Convergence for Nonconvex Optimization
In this section, we consider a nonconvex objective function. We establish global sublinear, local sublinear, and local linear convergence (see Theorem E.1 and Theorem E.3 for global and local sublinear convergence, respectively). We show that the FW algorithm converges to a minimizer $x^{*}$ when an initial point is close enough to $x^{*}$. We need weak convexity and local quadratic growth for local convergence.
@@
$f$ is $\rho$-weakly convex and has local $\mu$-quadratic growth with $\zeta>0$.
Assumption 5.1 is not too restrictive because the (local) strong convexity or the (local) Polyak–Łojasiewicz (PL) inequality implies the (local) quadratic growth condition for weakly convex functions (Liao et al., 2024, Theorem 3.1). For more examples satisfying Assumption 5.1 and $\rho\leq\mu$, see Examples E.4 and E.5.
Remark 5.2.
Linear convergence rates for nonconvex optimization problems are, to the best of our knowledge, the first such result for the Frank–Wolfe algorithm. Their proofs are technically challenging. Proposition C.8 is new and key for establishing a linear rate. In the nonconvex case, one cannot derive primal-gap inequalities such as Lemma D.1. We overcome this by restricting our function class to weakly convex functions and proving Lemma C.2. There remains an obstacle that could be resolved by the assumption of the quadratic growth condition. Assumption 5.1 enables us to derive Lemma C.3 and Proposition C.8 (as well as Remark C.4), which are crucial for establishing linear convergence.
We have local linear convergence, and its proof can be found in Appendix E.3.
@@
Finally, we establish the local linear convergence of Algorithm 3. Its proof can be found in Appendix E.4. We assume $\rho\leq L$, which is not restrictive (see Appendix E.2 and Example E.4).
@@
Without loss of generality, $\sigma=1$. If $\rho>L$, we can use $(\rho+L)D^{2}$, instead of $2LD^{2}$ as the initial bound of $f(x_{1})-f^{*}$. When $\phi=\frac{1}{2}\|\cdot\|^{2}$, i.e., $\nu=1$, Theorems 5.3 and 5.4 show linear convergence rates for Euclidean cases. Even if we use the Euclidean distance, our convergence results are novel in that linear convergence is achieved for the first time for a certain type of nonconvex problem. Moreover, when $P$ is $(\alpha,p)$-uniformly convex set (see Definition E.7) and $\phi=\frac{1}{2}\|\cdot\|^{2}$, we also establish local linear convergence (Theorems E.9 and E.10) in Appendix E.5.